# Coning off totally geodesic boundary components of a hyperbolic manifold

### v2

@@ 1. Introduction
The idea of coning off ends or boundary components of an aspherical manifold to obtain an aspherical space goes back at least to a comment of Gromov in [Gromov93, 7.A.VI]. This was fleshed out in the non-compact real hyperbolic case by Mosher and Sageev [MosherSageev]. Mosher and Sageev begin with a hyperbolic manifold with cusps. Under the condition that disjoint cusp cross sections with large (greater than $\pi$) injectivity radius can be chosen, they show that a negatively curved space can be obtained by cutting off the cusps and replacing them with cones. The method is differential geometric, utilizing the same kind of warped product construction as in the Gromov-Thurston $2\pi$ theorem [BleilerHodgson]. Modifying the construction, a wide variety of compact non-positively curved spaces can be obtained by replacing the cusps with mapping cylinders of fibrations [FM]. (In this paper non-positively curved means locally $\operatorname{CAT}(0)$, whereas negatively curved means locally $\operatorname{CAT}(k)$ for some $k<0$.)
@@
We also give conditions of the same flavor under which a closed locally convex subset of the original manifold gives a locally convex (and hence $\pi_{1}$–injective) subset of the coned-off space. These results are used in work of the second author with Ruffoni to give examples of negatively curved three-dimensional pseudo-manifolds whose fundamental groups do not act properly on any $\operatorname{CAT}(0)$ cube complex [MR].
As a consequence of our geometric construction, we obtain group-theoretic conclusions about the fundamental groups of the coned-off spaces and of the coned-off locally convex subsets. These are quantitative versions of results in “group-theoretic Dehn filling” [GM, Osin, GMQC].
@@ 1.1. Statement of main theorems
We describe the boundary at infinity of the universal cover of $\widehat{M}$ in Appendix A
@@
There are finitely many conjugacy classes $c_{1},\ldots,c_{k}$ in $\pi_{1}M_{0}$ representing nontrivial loops of length at most $\pi/\sinh(\operatorname{BW}_{M}(\partial M))$ in $\partial M_{0}$. Since $\pi_{1}M_{0}$ is linear, it is residually finite [Malcev40]. In particular there is a finite index normal subgroup of $\pi_{1}M_{0}$ missing the conjugacy classes $c_{1},\ldots,c_{k}$. The corresponding finite-sheeted cover $M\to M_{0}$ satisfies $\operatorname{inj}(\partial M)>\pi/\sinh(\operatorname{BW}_{M}(\partial M))$. In particular Theorem A applies to $M$ by Remark 1.5. ∎
@@
Scott’s characterization of subgroup separability [Scott78] means that any compact subset $K$ of the cover $\widetilde{M}_{0}$ of $M_{0}$ corresponding to $\pi_{1}S_{0}$ embeds in a finite-sheeted cover $M$ of $M_{0}$. The locally convex set $S_{0}$ lifts to a set $S=\widetilde{S}_{0}$ in $\widetilde{M}_{0}$ with infinite buffer width. Taking $K$ to be a closed cc–neighborhood of $\widetilde{S}_{0}$, the corresponding finite cover $M$ contains a lifted copy $S$ of $S_{0}$, which moreover satisfies condition 3. ∎
We remark that the condition of separability is always true if $\dim(M)\leq 3$ by results of Scott [Scott78] and Wise [Wise21].
@@
The first part of the Corollary is immediate from Theorem A. If each component of $\partial M$ intersects $S$ either in a connected set or in the empty set, the second part is immediate from Theorem B, since the image $\bar{H}$ of $H$ in $\Gamma/K$ can be identified with the fundamental group of $\widehat{S}$. Since $\widehat{M}$ is negatively curved and $\widehat{S}$ is a locally convex subset, the fundamental group of $\widehat{S}$ injects into that of $\widehat{M}$ (see [BH, II.4.14]). If there are components of $\partial M$ whose intersection with $S$ is disconnected, then $\pi_{1}\widehat{S}$ is still a quasi-convex subgroup, but is not equal to $\bar{H}$; rather, it contains $\bar{H}$ as a free factor by van Kampen’s theorem. But a free factor of a group is always quasi-isometrically embedded, so the quasi-convexity of $\bar{H}$ follows from the quasi-convexity of $\pi_{1}\widehat{S}$. ∎
The first part of Corollary 1.11 is related to the Relatively Hyperbolic Dehn Filling Theorem [GM, Osin] – the statement about cohomological dimension is related to work of Sun and Petrosyan [SP]. The second part is connected to [GMQC, Proposition 2.5]. All those statements though are for “sufficiently large” or “sufficiently HH–wide” fillings. In principle the bounds for those results can be made explicit, but will not be as tight as the ones which we have provided.
@@ 2.1. Warped products
We recall here the definition of a warped product of length spaces, which generalizes a warped product of Riemannian manifolds. For more information see [BBI, §3.6.4].
@@
Let $(X,d_{X})$ and $(Y,d_{Y})$ be length spaces and let $f\colon\thinspace X\to[0,\infty)$ be a continuous function. For a Lipschitz path $\gamma\colon\thinspace[a,b]\to X\times Y$, write $\gamma_{X}$ and $\gamma_{Y}$ for the projections to $X$ and $Y$ respectively. These are also Lipschitz paths so their speeds $|\gamma_{X}^{\prime}|$, $|\gamma_{Y}^{\prime}|$ are defined almost everywhere [BBI, Theorem 2.7.6]. Define a length
@@ 2.2. Local convexity
Pass to a finite subcover and let $\varepsilon$ be the minimum $r_{x}$ appearing in the subcover. We aim to show that $\mathrm{lcr}_{S}(S)\geq\frac{\varepsilon}{2}$. Fix some $B_{\frac{r_{x}}{2}}(x)$ in the finite subcover, and for readability write $\frac{r_{x}}{2}=r_{0}$. Let $c$ be a point in $B_{r_{0}}(x)$. We shall show $S\cap B_{r_{0}}(c)$ is convex in $Z$. Observe:
$S\cap B_{r_{0}}(c)$ $=S\cap B_{r_{0}}(c)\cap B_{r_{x}}(x)$
$=\left(B_{r_{0}}(c)\cap B_{r_{x}}(x)\right)\cap\left(S\cap B_{r_{x}}(x)\right).$
We claim that this is an intersection of two convex sets, hence convex. Indeed, $S\cap B_{\mathrm{lcr}_{S}(s)}(s)$ is convex by definition. As $B_{r_{x}}(x)$ contains $B_{r_{0}}(c)$ and $r_{x}<\operatorname{BW}_{Z}(x)$, we can lift $B_{r_{x}}(x)$ to the universal cover of $Z$, which is $\operatorname{CAT}(0)$, so balls are convex, showing that $B_{r_{0}}(c)\cap B_{r_{x}}(x)$ is convex. Therefore,
$\mathrm{lcr}_{S}(c)\geq r_{0}\geq\frac{\varepsilon}{2}.$
@@ 3. Tailoring warped products to bound curvature from above
Our first goal is to construct a warping function $f$ on an interval $I\subset\mathbb{R}$ such that $f$ vanishes at the left-hand endpoint of $I$, $f$ makes the metric on the cone agree with the hyperbolic metric on $M$ away from a neighborhood of the cone point, and the cone $I\times_{f}N$ is $\operatorname{CAT}(K)$ for some $K<0$, where $N$ is a boundary component of $M$. We will use a criterion of Alexander and Bishop [AB] to bound the curvature of our warped product from above. They work in a more general setting $B\times_{f}F$ where $B$ need not be an interval, so we simplify their definitions and result here.
@@
[AB, §2.2] Let $\mathcal{F}K$ denote the family of solutions to the differential equation $g^{\prime\prime}+Kg=0$. An $\mathcal{F}K$-convex function $f:I\to\mathbb{R}$ for $I\subseteq\mathbb{R}$ an interval is one whose restriction to every sub-interval satisfies the differential inequality $f^{\prime\prime}\geq-Kf$ in the barrier sense, meaning if $g\in\mathcal{F}K$ coincides with $f$ at the endpoints of a sub-interval, then $f\leq g$.
@@
[AB, Theorem 1.1] Let $I=[t_{0},t_{1})\subset\mathbb{R}$ and let $F$ be a complete $\operatorname{CAT}(K_{F})$ space for some $K_{F}\in\mathbb{R}$. Fix some $K\leq 0$ and suppose $f:I\to[0,\infty)$ is $\mathcal{F}K$-convex, with $f^{-1}(0)=\{t_{0}\}$. Suppose further that $f^{\prime}({t_{0}}^{+})^{2}\geq K_{F}$. [^3] Here $f^{\prime}(t_{0}^{+})$ denotes the right-hand derivative at $t_{0}$.
@@
Notice first that the hypotheses imply that $f$ is strictly increasing on $I$. For $t\in(t_{0},t_{1})$, let
@@
Lemma 2.3.(2) implies that $L_{t}$ is convex in $I\times_{f}F$. Moreover, [AB, Theorem 1.1] implies that $L_{t}$ is $\operatorname{CAT}(K)$. Any geodesic triangle lies in $L_{t}$ for some $t<t_{1}$, so the entire warped product is $\operatorname{CAT}(K)$. ∎
@@
Suppose by contradiction that $h(t_{0})>0$ for some $t_{0}\in J$. Shrinking $J$ if necessary, we may suppose that $h$ is positive everywhere on the interior of $J$. Almost everywhere on $J$, we have
@@
Choose $t_{0}<b$ so that the line of slope $\delta$ through the point $(t_{0},0)$ intersects $l_{0}$ somewhere on the interval $(t_{0},b)$. Let $l_{1}$ be this second line. By [FM, Lemma 4.13], there is a smooth function $g$ on $t_{0},b]$ whose graph is tangent to $l_{0}$ at $(b,\cosh(b))$ and to $l_{1}$ at $(t_{0},b)$, and whose second derivative is bounded below everywhere. In particular $g$ has the following properties, for some $\mu>0$.
@@
Since the sectional curvature of $N$ is bounded above by 0, so too is the sectional curvature of $\widetilde{N}$. Thus $\widetilde{N}$ is $\operatorname{CAT}(0)$ (cf. Theorem II.1A.6 in [BH]), which implies it is $\operatorname{CAT}(\kappa)$ (see [BH, Theorem II.1.12]). The ball $B$ lifts isometrically to $\widetilde{N}$, so $d_{B}(x,y)$ is bounded above by the corresponding distance in the $\operatorname{CAT}(\kappa)$ comparison triangle. By (2) this also gives an upper bound for $d_{N}(x,y)$, verifying the $\operatorname{CAT}(\kappa)$ inequality. ∎
@@ 5. Proof of Theorem B
(5) $\varepsilon<\min\left\{\frac{b^{\prime}-t(y)}{3},\frac{L}{2},\frac{cf(L)}{2},\frac{\mathrm{lcr}_{S}(S)}{2}\right\}.$
@@
Because $\varepsilon<\frac{1}{2}L=\frac{1}{3}(t(y)-t_{0})$, the ball $B_{\varepsilon}(y)$ is disjoint from $t_{0},L]\times N\subset C_{f}(N)$. Since the warping function $f$ is increasing and $B_{\varepsilon}(y)$ is entirely above $\{L\}\times N$,
@@
In [Sw20], Świątkowski defines the tree of manifolds $\mathcal{X}(\mathcal{M})$ based on a collection of manifolds $\mathcal{M}$ as a certain inverse limit of connect sums of elements of $\mathcal{M}$. We refer to [Sw20] for the relevant definitions. We show in this appendix that the boundary at infinity of the universal cover of $\widehat{M}$ is such a tree of manifolds. Note that this boundary is also the Gromov boundary of $\pi_{1}\widehat{M}$.
@@
To prove Theorem A.1 we will need some more information about the cones described in Lemma 4.1. Recall that the space of directions at $p$, written $\Sigma_{p}$, is a set of equivalence classes of (nondegenerate) geodesic segments starting at $p$. Two geodesic segments are equivalent if the Alexandrov angle between them is zero. The Alexandrov angle makes $\Sigma_{p}$ into a metric space (see [BH, II.3] for more detail). At any point of a Riemannian manifold, the space of directions can be identified with the unit tangent sphere, and the Alexandrov angle is the angle given by the Riemannian metric at that point. The cone described in Lemma 4.1 is Riemannian everywhere except for the cone point, where the space of directions is described by the following lemma.
@@
[AB] Let $N$ and $C_{f}(N)$ be as in Lemma 4.1. The space of directions at the cone point is homeomorphic to $N$ and Alexandrov angle in this space is given by
@@
We recall that $f\colon\thinspace[t_{0},\infty)\to[0,\infty)$ is a cone warping function with parameters $(b,\delta,K)$, as in Definition 3.5. In particular the function $f$ vanishes at $t_{0}$, and $f^{\prime}(t_{0}^{+})=\delta$. Recall also that $C_{f}(N)$ is a warped product with fiber $N$ and warping function $f$. This description of the space of directions at the cone point of such a warped product follows from [FM, Proposition 3.7], which is a restatement of results from [AB, p. 1147]. ∎
@@
We must verify that the universal cover of our model $X\cong\widehat{M}$ is a Riemannian $(\mathcal{N},0)$–pseudomanifold with log-injective singularities, in the sense of [Sw20, Definition 6.1]. We can then apply [Sw20, Theorem 6.2] to deduce that the visual boundary is homeomorphic to the tree of manifolds $X(\mathcal{N})$ as desired. The definition [Sw20, Definition 6.1] consists of six conditions. The first four conditions are immediate from the fact that our space $X$ is a compact nonpositively curved pseudomanifold with isolated singularities, whose metric is Riemannian away from those singularities.
@@
where $t_{1}-t_{0}$ is small enough so that $U$ is convex. Now suppose two points $x=(t,n_{1})$ and $y=(s,n_{2})$ in $U\smallsetminus\{p\}$ satisfy $\log_{p}(x)=\log_{p}(y)$. Since the second factor of the logarithm records the distance, we have $t=s$. By Lemma 2.3.(1), each of the geodesics $\sigma_{x}$ and $\sigma_{y}$ has constant second coordinate. Lemma A.2 tells us that the Alexandrov angle between $\sigma_{x}$ and $\sigma_{y}$ is the maximum of $\pi$ and $\frac{1}{\delta}d_{N}(n_{1},n_{2})$. In particular it can only be zero if $n_{1}=n_{2}$ and $x=y$.

### v3

@@ 1. Introduction
The idea of coning off ends or boundary components of an aspherical manifold to obtain an aspherical space goes back at least to a comment of Gromov in [6, 7.A.VI]. This was fleshed out in the non-compact real hyperbolic case by Mosher and Sageev [11]. Mosher and Sageev begin with a hyperbolic manifold with cusps. Under the condition that disjoint cusp cross sections with large (greater than $\pi$) injectivity radius can be chosen, they show that a negatively curved space can be obtained by cutting off the cusps and replacing them with cones. The method is differential geometric, utilizing the same kind of warped product construction as in the Gromov-Thurston $2\pi$ theorem [2]. Modifying the construction, a wide variety of compact non-positively curved spaces can be obtained by replacing the cusps with mapping cylinders of fibrations [5]. (In this paper non-positively curved means locally $\operatorname{CAT}(0)$, whereas negatively curved means locally $\operatorname{CAT}(k)$ for some $k<0$.)
@@
We also give conditions of the same flavor under which a closed locally convex subset of the original manifold gives a locally convex (and hence $\pi_{1}$–injective) subset of the coned-off space. These results are used in work of the second author with Ruffoni to give examples of negatively curved three-dimensional pseudo-manifolds whose fundamental groups do not act properly on any $\operatorname{CAT}(0)$ cube complex [10].
As a consequence of our geometric construction, we obtain group-theoretic conclusions about the fundamental groups of the coned-off spaces and of the coned-off locally convex subsets. These are quantitative versions of results in “group-theoretic Dehn filling” [7, 12, 8].
@@ 1.1. Statement of main theorems
We describe the boundary at infinity of the universal cover of $\widehat{M}$ in Appendix A.
@@
There are finitely many conjugacy classes $c_{1},\ldots,c_{k}$ in $\pi_{1}M_{0}$ representing nontrivial loops of length at most $\pi/\sinh(\operatorname{BW}_{M}(\partial M))$ in $\partial M_{0}$. Since $\pi_{1}M_{0}$ is linear, it is residually finite [9]. In particular there is a finite index normal subgroup of $\pi_{1}M_{0}$ missing the conjugacy classes $c_{1},\ldots,c_{k}$. The corresponding finite-sheeted cover $M\to M_{0}$ satisfies $\operatorname{inj}(\partial M)>\pi/\sinh(\operatorname{BW}_{M}(\partial M))$. In particular Theorem A applies to $M$ by Remark 1.5. ∎
@@
Scott’s characterization of subgroup separability [14] means that any compact subset $K$ of the cover $\widetilde{M}_{0}$ of $M_{0}$ corresponding to $\pi_{1}S_{0}$ embeds in a finite-sheeted cover $M$ of $M_{0}$. The locally convex set $S_{0}$ lifts to a set $S=\widetilde{S}_{0}$ in $\widetilde{M}_{0}$ with infinite buffer width. Taking $K$ to be a closed cc–neighborhood of $\widetilde{S}_{0}$, the corresponding finite cover $M$ contains a lifted copy $S$ of $S_{0}$, which moreover satisfies condition 3. ∎
We remark that the condition of separability is always true if $\dim(M)\leq 3$ by results of Scott [14] and Wise [16].
@@
The first part of the Corollary is immediate from Theorem A. If each component of $\partial M$ intersects $S$ either in a connected set or in the empty set, the second part is immediate from Theorem B, since the image $\bar{H}$ of $H$ in $\Gamma/K$ can be identified with the fundamental group of $\widehat{S}$. Since $\widehat{M}$ is negatively curved and $\widehat{S}$ is a locally convex subset, the fundamental group of $\widehat{S}$ injects into that of $\widehat{M}$ (see [3, II.4.14]). If there are components of $\partial M$ whose intersection with $S$ is disconnected, then $\pi_{1}\widehat{S}$ is still a quasi-convex subgroup, but is not equal to $\bar{H}$; rather, it contains $\bar{H}$ as a free factor by van Kampen’s theorem. But a free factor of a group is always quasi-isometrically embedded, so the quasi-convexity of $\bar{H}$ follows from the quasi-convexity of $\pi_{1}\widehat{S}$. ∎
The first part of Corollary 1.11 is related to the Relatively Hyperbolic Dehn Filling Theorem [7, 12] – the statement about cohomological dimension is related to work of Sun and Petrosyan [13]. The second part is connected to [8, Proposition 2.5]. All those statements though are for “sufficiently large” or “sufficiently HH–wide” fillings. In principle the bounds for those results can be made explicit, but will not be as tight as the ones which we have provided.
@@ 2.1. Warped products
We recall here the definition of a warped product of length spaces, which generalizes a warped product of Riemannian manifolds. For more information see [4, §3.6.4].
@@
Let $(X,d_{X})$ and $(Y,d_{Y})$ be length spaces and let $f\colon\thinspace X\to[0,\infty)$ be a continuous function. For a Lipschitz path $\gamma\colon\thinspace[a,b]\to X\times Y$, write $\gamma_{X}$ and $\gamma_{Y}$ for the projections to $X$ and $Y$ respectively. These are also Lipschitz paths so their speeds $|\gamma_{X}^{\prime}|$, $|\gamma_{Y}^{\prime}|$ are defined almost everywhere [4, Theorem 2.7.6]. Define a length
@@ 2.2. Local convexity
By compactness there is a finite set $\{x_{1},\ldots,x_{m}\}$ so that for every $c\in S$ there is an $i$ so that $c$ is contained in the open $r_{x_{i}}/2$–ball around $x_{i}$. We set $r_{i}=r_{x_{i}}$. Let $\varepsilon=\min\{r_{1},\ldots,r_{m}\}$. We will show that $\mathrm{lcr}_{S}(S)\geq\frac{\varepsilon}{2}$.
Fix $c\in S$ and let $x=x_{i}$ for some $i$ with $d_{Z}(c,x_{i})<\frac{r_{i}}{2}$. Set $\delta=\frac{r_{i}}{2}$, and note $\delta\geq\frac{\varepsilon}{2}$, so it suffices to show that $S\cap B_{\delta}(c)$ is convex in $Z$. Observe:
$S\cap B_{\delta}(c)$ $=S\cap B_{r_{x}}(x)\cap B_{\delta}(c)$
$=\bigl(S\cap B_{r_{x}}(x)\bigr)\cap\bigl(B_{\delta}(c)\cap B_{r_{x}}(x)\bigr).$
We claim that this is an intersection of two convex sets, and hence convex. Indeed, $S\cap B_{r_{x}}(x)$ is convex by definition. As $B_{r_{x}}(x)$ contains $B_{\delta}(c)$ and $r_{x}<\operatorname{BW}_{Z}(x)$, we can lift $B_{r_{x}}(x)$ to the universal cover of $Z$, where balls are convex by the nonpositive curvature condition. Thus $B_{\delta}(c)\cap B_{r_{x}}(x)=B_{\delta}(c)$ is convex, and therefore so is $S\cap B_{\delta}(c)$. ∎
@@ 3. Tailoring warped products to bound curvature from above
Our first goal is to construct a warping function $f$ on an interval $I\subset\mathbb{R}$ such that $f$ vanishes at the left-hand endpoint of $I$, $f$ makes the metric on the cone agree with the hyperbolic metric on $M$ away from a neighborhood of the cone point, and the cone $I\times_{f}N$ is $\operatorname{CAT}(K)$ for some $K<0$, where $N$ is a boundary component of $M$. We will use a criterion of Alexander and Bishop [1] to bound the curvature of our warped product from above. They work in a more general setting $B\times_{f}F$ where $B$ need not be an interval, so we simplify their definitions and result here.
@@
[1, §2.2] Let $\mathcal{F}K$ denote the family of solutions to the differential equation $g^{\prime\prime}+Kg=0$. An $\mathcal{F}K$-convex function $f:I\to\mathbb{R}$ for $I\subseteq\mathbb{R}$ an interval is one whose restriction to every sub-interval satisfies the differential inequality $f^{\prime\prime}\geq-Kf$ in the barrier sense, meaning if $g\in\mathcal{F}K$ coincides with $f$ at the endpoints of a sub-interval, then $f\leq g$.
We remark that for $K\leq 0$, an $\mathcal{F}K$-convex function is convex in the usual sense.
@@
[1, Theorem 1.1] Let $I=[t_{0},t_{1})\subset\mathbb{R}$ and let $F$ be a complete $\operatorname{CAT}(K_{F})$ space for some $K_{F}\in\mathbb{R}$. Fix some $K\leq 0$ and suppose $f:I\to[0,\infty)$ is $\mathcal{F}K$-convex, with $f^{-1}(0)=\{t_{0}\}$. Suppose further that $f^{\prime}({t_{0}}^{+})^{2}\geq K_{F}$. [^3] Here $f^{\prime}(t_{0}^{+})$ denotes the right-hand derivative at $t_{0}$.
@@
Notice first that since $f$ is convex, non-negative and satisfies $f^{-1}(0)=\{t_{0}\}$, it is strictly increasing on $I$. For $t\in(t_{0},t_{1})$, let
@@
Lemma 2.3.(2) implies that $L_{t}$ is convex in $I\times_{f}F$. Moreover, [1, Theorem 1.1] implies that $L_{t}$ is $\operatorname{CAT}(K)$. Any geodesic triangle in $I\times_{f}F$ lies in $L_{t}$ for some $t<t_{1}$, so the entire warped product is $\operatorname{CAT}(K)$. ∎
@@
Suppose by contradiction that $h(t_{0})>0$ for some $t_{0}\in J$. Shrinking $J$ if necessary, we may suppose that $h$ is positive everywhere on the interior of $J$ and vanishes at the endpoints of $J$. Almost everywhere on $J$, we have
@@
Choose $t_{0}<b$ so that the line of slope $\delta$ through the point $(t_{0},0)$ intersects $l_{0}$ somewhere on the interval $(t_{0},b)$. Let $l_{1}$ be this second line. By [5, Lemma 4.13], there is a smooth function $g$ on $t_{0},b]$ whose graph is tangent to $l_{0}$ at $(b,\cosh(b))$ and to $l_{1}$ at $(t_{0},0)$, and whose second derivative is bounded below everywhere. In particular $g$ has the following properties, for some $\mu>0$.
@@
Since the sectional curvature of $N$ is bounded above by 0, so too is the sectional curvature of $\widetilde{N}$. Thus $\widetilde{N}$ is $\operatorname{CAT}(0)$ (cf. Theorem II.1A.6 in [3]), which implies it is $\operatorname{CAT}(\kappa)$ (see [3, Theorem II.1.12]). The ball $B$ lifts isometrically to $\widetilde{N}$, so $d_{B}(x,y)$ is bounded above by the corresponding distance in the $\operatorname{CAT}(\kappa)$ comparison triangle. By (2) this also gives an upper bound for $d_{N}(x,y)$, verifying the $\operatorname{CAT}(\kappa)$ inequality. ∎
@@ 5. Proof of Theorem B
(5) $\varepsilon<\min\left\{\frac{b^{\prime}-t(y)}{3},\frac{t(y)-L}{2},\frac{cf(L)}{2},\frac{\mathrm{lcr}_{S}(S)}{2}\right\}.$
@@
Because $\varepsilon<\frac{1}{2}(t(y)-L)$, the ball $B_{\varepsilon}(y)$ is disjoint from $t_{0},L]\times N\subset C_{f}(N)$. Since the warping function $f$ is increasing and $B_{\varepsilon}(y)$ is entirely above $\{L\}\times N$,
@@
In [15], Świątkowski defines the tree of manifolds $\mathcal{X}(\mathcal{M})$ based on a collection of manifolds $\mathcal{M}$ as a certain inverse limit of connected sums of elements of $\mathcal{M}$. We refer to [15] for the relevant definitions. We show in this appendix that the boundary at infinity of the universal cover of $\widehat{M}$ is such a tree of manifolds. Note that this boundary is also the Gromov boundary of $\pi_{1}\widehat{M}$.
@@
To prove Theorem A.1 we will need some more information about the cones described in Lemma 4.1. Recall that the space of directions at $p$, written $\Sigma_{p}$, is a set of equivalence classes of (nondegenerate) geodesic segments starting at $p$. Two geodesic segments are equivalent if the Alexandrov angle between them is zero. The Alexandrov angle makes $\Sigma_{p}$ into a metric space (see [3, II.3] for more detail). At any point of a Riemannian manifold, the space of directions can be identified with the unit tangent sphere, and the Alexandrov angle is the angle given by the Riemannian metric at that point. The cone described in Lemma 4.1 is Riemannian everywhere except for the cone point, where the space of directions is described by the following lemma.
@@
[1] Let $N$ and $C_{f}(N)$ be as in Lemma 4.1. The space of directions at the cone point is homeomorphic to $N$ and Alexandrov angle in this space is given by
@@
We recall that $f\colon\thinspace[t_{0},\infty)\to[0,\infty)$ is a cone warping function with parameters $(b,\delta,K)$, as in Definition 3.5. In particular the function $f$ vanishes at $t_{0}$, and $f^{\prime}(t_{0}^{+})=\delta$. Recall also that $C_{f}(N)$ is a warped product with fiber $N$ and warping function $f$. This description of the space of directions at the cone point of such a warped product follows from [5, Proposition 3.7], which is a restatement of results from [1, p. 1147]. ∎
@@
We must verify that the universal cover of our model $X\cong\widehat{M}$ is a Riemannian $(\mathcal{N},0)$–pseudomanifold with log-injective singularities, in the sense of [15, Definition 6.1]. We can then apply [15, Theorem 6.2] to deduce that the visual boundary is homeomorphic to the tree of manifolds $X(\mathcal{N})$ as desired. The definition [15, Definition 6.1] consists of six conditions. The first four conditions are immediate from the fact that our space $X$ is a compact nonpositively curved pseudomanifold with isolated singularities, whose metric is Riemannian away from those singularities.
@@
where $t_{1}-t_{0}$ is small enough so that $U$ is convex. Now suppose two points $x=(t,n_{1})$ and $y=(s,n_{2})$ in $U\smallsetminus\{p\}$ satisfy $\log_{p}(x)=\log_{p}(y)$. Since the second factor of the logarithm records the distance, we have $t=s$. By Lemma 2.3.(1), each of the geodesics $\sigma_{x}$ and $\sigma_{y}$ has constant second coordinate. Lemma A.2 tells us that the Alexandrov angle between $\sigma_{x}$ and $\sigma_{y}$ is the minimum of $\pi$ and $\frac{1}{\delta}d_{N}(n_{1},n_{2})$. In particular it can only be zero if $n_{1}=n_{2}$ and $x=y$.